# Monotone Subsystem Decomposition for Efficient Multi-Objective Robot Design

### v1

@@ III-A Single-Objective Constrained Optimization
Constraints are mathematical expressions that describe relationships between the properties of components. These constraints can not only specify valid component combinations but also describe the performance specifications of the robot. Inequality constraints take the general form of $L_{c}(D_{v,p},D_{v^{\prime},p^{\prime}},...)\leq R_{c}(D_{v,p},D_{v^{\prime},p^{\prime}},...)$ where $L_{c}:\mathbb{R}^{\sum_{v}|P_{v}|}\rightarrow\mathbb{R}$ and $R_{c}:\mathbb{R}^{\sum_{v}|P_{v}|}\rightarrow\mathbb{R}$ are real-valued functions that operate on the variable properties and describe constraint $c\in C$.
@@
$4*(M_{\textrm{voltage}}*M_{\textrm{current}})\vspace{-1.25mm}\leq B_{\textrm{voltage}}*B_{\textrm{current}}$
@@ IV PROBLEM FORMULATION
We can determine a subsystem ${X}$ within the constraint graph by selecting a subset of variables from the larger system ${Z}={X}\cup{Y}$, where ${X}\subset{Z}\subseteq{V}$ and ${Y}={Z}\setminus{X}$. We denote an implementation $\mathbf{z}$ of subsystem ${Z}$ that uses implementations $\mathbf{x}$ and $\mathbf{y}$ from ${X},{Y}$ as $\mathbf{z}=\mathbf{x}\cup\mathbf{y}$. A system ${X}$ shares a constraint or objective function with a system ${Y}$ if a variable property from both ${X}$ and ${Y}$ is included in the domain of that constraint or function. We call a subsystem fully consistent if all constraint and objective functions that are shared between ${X}$ and ${Y}$ are fully consistent.
@@ V THEORETICAL RESULTS
Theorem: Consider the optimization of a system ${Z}={X}\cup{Y}$ where ${Z}\subseteq{V}$ and ${X}$ is a fully consistent subsystem within ${Z}$. Then all assignments of ${Z}$ on the Pareto front $O_{Z}$ can use assignments from the Pareto front $O_{X}$ of the subsystem ${X}$, where $O_{X}$ is optimized for variable properties used in shared functions between ${X}$ and ${V}\setminus{X}$.
@@
Proof: As ${X}$ is a fully consistent subsystem, its shared properties are either maximization or minimization properties.
Consider the maximization property $D^{\uparrow}_{x,p}\in D_{x}$ in $X$. Let $\mathbf{x^{\prime}}$ and $\mathbf{x}$ be assignments such that $\mathbf{x^{\prime}}\notin O_{X}$, ${\mathbf{x}}\in O_{X}$. Thus, $d^{\uparrow}_{\mathbf{x^{\prime}},p}\leq d^{\uparrow}_{\mathbf{{x}},p}$ for $d^{\uparrow}_{\mathbf{x^{\prime}},p},d^{\uparrow}_{\mathbf{{x}},p}\in D^{\uparrow}_{x,p}$ as $O_{X}$ was maximized for $D^{\uparrow}_{x,p}$. Additionally, let $f^{\uparrow}_{i}(D_{{x},p},D_{y,p^{\prime}},\ldots)$ and $f_{i}^{\downarrow}(D_{{x},p},D_{y,p^{\prime}},\ldots)$ represent objective functions in the system ${Z}$, and let $R_{c}(D_{x,p},D_{y,p^{\prime}},...)$ and $L_{c}(D_{x,p},D_{y,p^{\prime}},...)$ represent shared constraint functions in ${Z}$.
@@
Feasibility: Since the constraints are structured $L_{c}(D_{x,p},D_{y,p^{\prime}},...)\leq R_{c}(D_{x,p},D_{y,p^{\prime}},...)$, then $\mathbf{x}$ will not violate any constraints that $\mathbf{x^{\prime}}$ does not already violate since $L_{c}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},...)\leq L_{c}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},...)$ and $R_{c}(d^{\uparrow}_{\mathbf{x^{\prime}},p},d_{\mathbf{y},p^{\prime}},...)\leq R_{c}(d^{\uparrow}_{\mathbf{x},p},d_{\mathbf{y},p^{\prime}},...)$ as $R_{c}$ and $L_{c}$ are fully consistent.
@@ VII-A Quadcopter System Design Problem
Figure 3 compares the scaling performance of our CP approach against the LP approach. While the original quadcopter design problem is single-objective, our method extends to multi-objective optimization so we also report timing results for solving a multi-objective variant that minimizes system mass alongside maximizing velocity. As shown, our CP approach produces solutions faster than LP for all but the smallest of catalogs (in which solve times are less than a second). For catalogs exceeding 200 components, our CP approach even solves the multi-objective design problem and determines a PF of optimal solutions faster than the LP approach can solve the single-objective case. Our approach has a log-log slope $<1$, indicating its scaling is sub-linear.

### v2

@@ III-A Single-Objective Constrained Optimization
[…]
@@
$4*(M_{\textrm{voltage}}*M_{\textrm{current}})\vskip-3.55658pt\leq B_{\textrm{voltage}}*B_{\textrm{current}}$
@@ IV PROBLEM FORMULATION
We can determine a subsystem X{X} within the constraint graph by selecting a subset of variables from the larger system ${Z}={X}\cup{Y}$, where ${X}\subset{Z}\subseteq{V}$ and ${Y}={Z}\setminus{X}$. We denote an implementation $\mathbf{z}$ of subsystem Z{Z} that uses implementations $\mathbf{x}$ and $\mathbf{y}$ from X,Y{X},{Y} as $\mathbf{z}=\mathbf{x}\cup\mathbf{y}$. A system X{X} shares a constraint or objective function with a system Y{Y} if a variable property from both X{X} and Y{Y} is included in the domain of that constraint or function. We call a subsystem fully consistent if all constraint and objective functions that are shared between X{X} and Y{Y} are fully consistent.
@@ V THEORETICAL RESULTS
Theorem: Consider the optimization of a system ${Z}={X}\cup{Y}$ where ${Z}\subseteq{V}$ and X{X} is a fully consistent subsystem within Z{Z}. Then all assignments of Z{Z} on the Pareto front $O_{Z}$ can use assignments from the Pareto front $O_{X}$ of the subsystem X{X}, where $O_{X}$ is optimized for variable properties used in shared functions between X{X} and ${V}\setminus{X}$.
@@
Proof: As X{X} is a fully consistent subsystem, its shared properties are either maximization or minimization properties.
[…]
@@
[…]
@@ VII-A Quadcopter System Design Problem
Figure 3 compares the scaling performance of our CP approach against the LP approach. While the original quadcopter design problem is single-objective, our method extends to multi-objective optimization so we also report timing results for solving a multi-objective variant that minimizes system mass alongside maximizing velocity. As shown, our CP approach produces solutions faster than LP for all but the smallest of catalogs (in which solve times are less than a second). For catalogs exceeding 200 components, our CP approach even solves the multi-objective design problem and determines a PF of optimal solutions faster than the LP approach can solve the single-objective case. Our approach has a log-log slope <1<1, indicating its scaling is sub-linear.